# RBI-ThPhys-2024-15

### v2

@@ VI.2 Including higher-order corrections
where $\sigma_{\mathrm{LO}}$ is also computed for single-Higgs production for consistency using MadGraph5_amc@nlo. To define K-factors for the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms we propose the ansatz:
@@
As expected, Figures 1113 show a significant difference between the LO and NLO yields due to the differences in the cross-sections. However, after application of the K-factors the yields agree (by construction) and we observe a good agreement in the shapes of the distributions compared to the NLO, indicating that LO samples are adequate for describing the $m_{hh}$ spectrum. The shapes of the NLO-approx distributions also agree well with the NLO distributions, and in this case the yields/cross-section before application of the K-factors are also noticeable quite similar to the NLO results. Figures 15 and 16 show the LO and NLO-approx distributions for the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms, respectively. While no comparisons to the full NLO results are possible in this case, the agreement between the shapes and the yields (after K-factor scaling) of the LO and NLO-approx distributions gives further confidence in the description of the $m_{hh}$ spectrum by the LO MC samples after reweighting.
@@ VII.1 BM1
The interference tends increase (decrease) the broader non-resonant-like continuum above (below) the resonant peak at $m_{H}\sim 620\,{\,\rm GeV}$. After detector smearing, the overall distribution exhibits an interesting double-peak structure. The interference also causes a shift in the position of the mass peak to slightly lower values, however, after smearing the shift in the peak position is not very noticeable. It is clear from comparing to the $\mathrm{S_{H}}$-only spectrum that neglecting the non-resonant terms altogether results in a failure to describe the double-peak structure, and it also results in an underestimation of the number of events close to the peak of around 35%. Including the non-resonant terms while neglecting only the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms clearly brings an improvement in the overall description of the spectrum. However, it still results in an underestimate of the height of the $\mathrm{S_{H}}$ peak by about 15%, and an under(over)-estimate of the non-resonant continuum of $\approx$ 15–30% (15–40%) below (above) the resonant peak. This clearly motivates the proper consideration of the interference effects when performing resonant searches. We also do not observe a significant difference between no-interference and SM+$\mathrm{S_{H}}$ distributions. This is because only small values of $\sin\alpha$ are still allowed by the experimental constraints and this results in $\kappa_{\lambda_{hhh}}$ and $\kappa_{q}^{h}$ values that are typically close to unity at LO. As will be shown in the subsequent sections, this feature is common to all BM points.
@@ VII.5 BM8
The $m_{hh}$ distributions are shown in Figure 36. We observe no major differences between the no-interference and SM+$\mathrm{S_{H}}$ distributions, implying that assuming SM couplings for the $h$ is a reasonable assumption in this case. The non-resonant contribution to the spectrum is less noticeable in these plots as the resonant $\mathrm{S_{H}}$ peak is large. To study in more detail how the values of the $h$ couplings influence the non-resonant contribution, we remove the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ contributions to the spectrum, as shown in Figure 37. The non-resonant spectrum for BM8 shown in blue is compared to the SM shown in red. We see only small differences between the two distributions, mainly impacting the low $m_{hh}$ bins where the event yield is low. The differences in the distributions are somewhat smaller than one might expect given that $\kappa_{\lambda_{hhh}}$ deviates from the SM value by a non-negligible amount. This is due to a small deviation in the value of $\kappa_{q}^{h}$, which is equal to 0.97 in this BM. This causes an accidental cancellation between the $\Box$ and $\mathrm{S_{h}}$ components, which decrease in size for couplings $<1$, and the $-\mathrm{S_{h}}{\text{-}}\Box$ component which increases due to the destructive nature of the interference. The scenario where only $\kappa_{\lambda_{hhh}}$ deviates from the SM and $\kappa_{q}^{h}=1$ is shown in green in the Figure for comparison. Even in this scenario the deviation in the non-resonant yield is only $\approx 10\%$ so it is unlikely that the LHC experiments will be sensitive to the deviations in $\kappa_{\lambda_{hhh}}$ in this model [^18] We note that BM8 has since been excluded by the new combination presented by CMS in Ref. CMS:2024phk that is not yet included in HiggsTools. However, as the purpose of this benchmark is to study the most extreme cases where $\kappa_{\lambda_{hhh}}$ deviates from the SM, we still consider this a useful scenario since the new bounds from CMS restrict $\kappa_{\lambda_{hhh}}$ to be even closer to unity, and thus the conclusions drawn from this BM are still valid. We thank H. Bahl for useful discussions regarding this point. .
@@ VII.6 BM9
Figure 39: The di-Higgs mass distributions for BM9 after experimental smearing are displayed for the $s$-channel only ($\mathrm{S_{H}}$) contributions. The blue lines show distributions obtained using the correct width ($\Gamma_{H}=9.8\,{\,\rm GeV}$), while the red lines display the distributions obtained using the NWA with $\Gamma_{H}=20\,{\rm MeV}$ and rescaling the the branching fractions by $(20\,{\rm MeV})/\Gamma_{H}$ as described in the text. The left plot shows the distributions for the nominal assumed experimental resolution while the right plot shows the distributions after artificially improving the $m_{hh}$ resolution by a factor of two.
@@ VIII Summary
We have identified regions of the parameter space where interference effects between the non-resonant and resonant diagrams are sizeable. In order to account for NLO kinematics as well as interference effects, we have proposed a prescription that includes both, where we find good agreement with the correct NLO description when available. Using our methodology, we have defined benchmark scenarios exhibiting interesting features of the model. For several benchmarks scenarios, we find that the contribution of the non-resonant di-Higgs diagrams is significant and in some cases even dwarfs the resonant component. We have also observed that interference effects have a non-negligible effect on both the cross-sections and differential distributions. We strongly encourage the experimental collaborations to take these effects into account in future searches for additional Higgs bosons. Furthermore, the correct physical prediction of the total width needs to be used in realistic simulation as the width can crucially impact the rates as well as the relevance of the interference term. Finally, we have presented a publicly available tool, HHReweighter reweight_repo , for performing matrix-element reweighting that allows such interference effects to be taken into account in a computationally efficient way.
@@ Acknowledgements
TR is supported by the Croatian Science Foundation (HRZZ) under project IP-2022-10-2520. EF and FF acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967. DW was supported by the Science and Technology Facilities Council [grant number ST/W000636/1]. TR and EF also want to thank the CERN-TH group for their hospitality during parts of this work.

### v3

@@ VI.2 Including higher-order corrections
where $\sigma_{\mathrm{LO}}$ is also computed for single-Higgs production for consistency using MadGraph5. To define K-factors for the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms we propose the ansatz:
@@
As expected, Figures 1113 show a significant difference between the LO and NLO yields due to the differences in the cross-sections. However, after application of the K-factors the yields agree (by construction) and we observe a good agreement in the shapes of the distributions compared to the NLO, indicating that LO samples are adequate for describing the $m_{hh}$ spectrum. The shapes of the NLO-approx distributions also agree well with the NLO distributions, and in this case the yields/cross-section before application of the K-factors are also noticeably quite similar to the NLO results. We note that the agreement of the NLO-approx results with the NLO are slightly better for the $\Box$ contribution compared to the $\mathrm{S_{h}}$ contribution. This is because we are reweighting a SM di-Higgs sample where the $\Box$ diagram contributes more strongly to the total ME than the $\mathrm{S_{h}}$ and hence the weights used to obtain the $\Box$ contribution will tend to be closer to unity and thus less sensitive to the approximations of the method.
Figures 15 and 16 show the LO and NLO-approx distributions for the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms, respectively. While no comparisons to the full NLO results are possible in this case, the agreement between the shapes and the yields (after K-factor scaling) of the LO and NLO-approx distributions gives further confidence in the description of the $m_{hh}$ spectrum by the LO MC samples after reweighting.
@@ VII.1 BM1
We note that most of the current experimental searches neglect the non-resonant contributions to the spectrum entirely, which corresponds to the green line in Figure 18 and equivalent figures that will follow. However, a small number of searches fix the non-resonant di-Higgs process to the SM expectation and neglect the interference effects, which corresponds to the magenta line in the figures.
The interference tends to increase (decrease) the broader non-resonant-like continuum above (below) the resonant peak at $m_{H}\sim 620\,{\,\rm GeV}$. After detector smearing, the overall distribution exhibits an interesting double-peak structure. The interference also causes a shift in the position of the mass peak to slightly lower values, however, after smearing the shift in the peak position is not very noticeable. It is clear from comparing to the $\mathrm{S_{H}}$-only spectrum that neglecting the non-resonant terms altogether results in a failure to describe the double-peak structure, and it also results in an underestimation of the number of events close to the peak of around 35%. Including the non-resonant terms while neglecting only the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ interference terms clearly brings an improvement in the overall description of the spectrum. However, it still results in an underestimate of the height of the $\mathrm{S_{H}}$ peak by about 15%, and an under(over)-estimate of the non-resonant continuum of $\approx$ 15–30% (15–40%) below (above) the resonant peak. This clearly motivates the proper consideration of the interference effects when performing resonant searches. We also do not observe a significant difference between no-interference and SM+$\mathrm{S_{H}}$ distributions. This is because only small values of $\sin\alpha$ are still allowed by the experimental constraints and this results in $\kappa_{\lambda_{hhh}}$ and $\kappa_{q}^{h}$ values that are typically close to unity at LO. As will be shown in the subsequent sections, this feature is common to all BM points.
@@ VII.5 BM8
The $m_{hh}$ distributions are shown in Figure 36. We observe no major differences between the no-interference and SM+$\mathrm{S_{H}}$ distributions, implying that assuming SM couplings for the $h$ is a reasonable assumption in this case. We note that this conclusion does not imply that the interference effects are insignificant; it only indicates that fixing the $h$ couplings to the SM-values has a minimal impact on the distributions. The non-resonant contribution to the spectrum is less noticeable in these plots as the resonant $\mathrm{S_{H}}$ peak is large. To study in more detail how the values of the $h$ couplings influence the non-resonant contribution, we remove the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ contributions to the spectrum, as shown in Figure 37. The non-resonant spectrum for BM8 shown in blue is compared to the SM shown in red. We see only small differences between the two distributions, mainly impacting the low $m_{hh}$ bins where the event yield is low. The differences in the distributions are somewhat smaller than one might expect given that $\kappa_{\lambda_{hhh}}$ deviates from the SM value by a non-negligible amount. This is due to a small deviation in the value of $\kappa_{q}^{h}$, which is equal to 0.97 in this BM. This causes an accidental cancellation between the $\Box$ and $\mathrm{S_{h}}$ components, which decrease in size for couplings $<1$, and the $-\mathrm{S_{h}}{\text{-}}\Box$ component which increases due to the destructive nature of the interference. The scenario where only $\kappa_{\lambda_{hhh}}$ deviates from the SM and $\kappa_{q}^{h}=1$ is shown in green in Fig. 37 for comparison. Even in this scenario the deviation in the non-resonant yield is only $\approx 10\%$ so it is unlikely that the LHC experiments will be sensitive to the deviations in $\kappa_{\lambda_{hhh}}$ in this model [^19] We note that BM8 has since been excluded by the new combination presented by CMS in Ref. CMS:2024phk that is not yet included in HiggsTools. However, as the purpose of this benchmark is to study the most extreme cases where $\kappa_{\lambda_{hhh}}$ deviates from the SM, we still consider this a useful scenario since the new bounds from CMS restrict $\kappa_{\lambda_{hhh}}$ to be even closer to unity, and thus the conclusions drawn from this BM are still valid. We thank H. Bahl for useful discussions regarding this point. .
@@ VII.6 BM9
Figure 39: The di-Higgs mass distributions for BM9 after experimental smearing are displayed for the $s$-channel only ($\mathrm{S_{H}}$) contributions. The blue lines show distributions obtained using the correct width ($\Gamma_{H}=9.8\,{\,\rm GeV}$), while the red lines display the distributions obtained using the NWA with $\Gamma_{H}=20\,{\rm MeV}$ and rescaling the branching fractions by $(20\,{\rm MeV})/\Gamma_{H}$ as described in the text. The left plot shows the distributions for the nominal assumed experimental resolution while the right plot shows the distributions after artificially improving the $m_{hh}$ resolution by a factor of two.
@@ VIII Summary
We have identified regions of the parameter space where interference effects between the non-resonant and resonant diagrams are sizeable. In order to account for NLO kinematics as well as interference effects, we have proposed a prescription that includes both, where we find good agreement with the correct NLO description when available. Using our methodology, we have defined benchmark scenarios exhibiting interesting features of the model. For several benchmarks scenarios, we find that the contribution of the non-resonant di-Higgs diagrams is significant and in some cases even dwarfs the resonant component. We have also observed that interference effects have a non-negligible effect on both the cross-sections and differential distributions. We strongly encourage the experimental collaborations to take these effects into account in future searches for additional Higgs bosons. Furthermore, the correct physical prediction of the total width needs to be used in realistic simulation as the width can crucially impact the rates as well as the relevance of the interference term. We have presented a publicly available tool, HHReweighter reweight_repo , for performing matrix-element reweighting that allows such interference effects to be taken into account in a computationally efficient way.
Finally, we want to point out that the model studied here is already vastly explored by the experimental collaborations, and very often limits are set neglecting the effects discussed here. Therefore, even for such simple scenarios, results as currently presented should be taken with a grain of salt and in fact be reinterpreted using e.g. our reweighting tool. It would also be desirable to have experimental results presented e.g. in multi-dimensional grids, the simplest being additional results for varying widths for a given new scalar mass. Of course for other new physics scenarios the respective parameter space would have to be extended.
@@ Acknowledgements
TR is supported by the Croatian Science Foundation (HRZZ) under project HRZZ-IP-2022-10-2520. EF and FF acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967. DW was supported by the Science and Technology Facilities Council [grant number ST/W000636/1]. TR and EF also want to thank the CERN-TH group for their hospitality during parts of this work.